If a 64 foot rope weighs 20 lbs how much does an 80 foot rope weigh?
Answer by jim_thompson5910(35256)

64/20 = 80/x


64x = 20*80


64x = 1600


x = 1600/64


x = 25


So an 80 ft rope weighs 25 pounds

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim